### v1

# Diffusion Self-Weighted Guidance for Offline Reinforcement Learning

Abstract
Offline reinforcement learning (RL) recovers the optimal policy $\pi$ given historical observations of an agent. In practice, $\pi$ is modeled as a weighted version of the agent’s behavior policy $\mu$, using a weight function $w$ working as a critic of the agent’s behavior. Though recent approaches to offline RL based on diffusion models have exhibited promising results, the computation of the required scores is challenging due to their dependence on the unknown $w$. In this work, we alleviate this issue by constructing a diffusion over both the actions and the weights. With the proposed setting, the required scores are directly obtained from the diffusion model without learning extra networks. Our main conceptual contribution is a novel guidance method, where guidance (which is a function of $w$) comes from the same diffusion model, therefore, our proposal is termed Self-Weighted Guidance (SWG). We show that SWG generates samples from the desired distribution on toy examples and performs on par with state-of-the-art methods on D4RL’s challenging environments, while maintaining a streamlined training pipeline. We further validate SWG through ablation studies on weight formulations and scalability.
1 Introduction
Offline reinforcement learning (RL) learns a policy from a dataset of an agent’s collected interactions. This setting is commonplace in scenarios such as robotics and healthcare, where real-world exploration with suboptimal agents is costly, dangerous, and thus unfeasible levine2020offlinereinforcementlearningtutorial .
A critical issue when estimating the state-action value function $Q(s,\mathbf{a})$ (Q-values) in offline RL is the unknown effects of actions absent in the dataset, referred to out-of-sample actions. In practice, methods such as iql ; offlinerl_no_ood ; idql refrain from querying out-of-sample actions during training, thus ensuring that the learned policy remains close to the dataset policy, and avoiding distributional shift bear ; bcq ; nair2020awac . These methods typically adopt a weighted behavior policy form given by
$\pi(\mathbf{a}|s)\propto\mu(\mathbf{a}|s)w(s,\mathbf{a}),$ (1)
where $(s,\mathbf{a})$ is the state-action pair, $\pi(\mathbf{a}|s)$ is the target policy, $\mu(\mathbf{a}|s)$ is the dataset —or behavior— policy, and $w(s,\mathbf{a})$ is the weight function encoding the value of an action given a state.
Previous approaches learn $\pi(\mathbf{a}|s)$ in a weighted regression procedure weighted_regression ; AWRPeng19 ; nair2020awac ; critic_regularized_regression , using, e.g., a Gaussian model iql ; this approach is severely limited by the expressiveness of the chosen model. For instance, when the dataset contains multi-modal actions, uni-modal policies will certainly fail to capture the agent’s behavior successfully idql ; diff_dice .
Diffusion models (DM) sohl2015deep ; ddpm ; songscore have been successfully applied to sampling from the unnormalized target policy in Eq. (1) janner2022diffuser ; sfbc ; idql ; cep ; diff_dice , while modeling a potentially multi-modal policy dql ; diffusionpolicy . This has been achieved by approximating the behavior policy using a DM, i.e., $\mu_{\theta}(\mathbf{a}|s)\approx\mu(\mathbf{a}|s)$, learning the weight function using external models (see Sec. 2.1 and 2.2), and then constructing a procedure to emulate samples from $\pi(\mathbf{a}|s)$. The main challenge of this approach is the estimation of the intractable score, where the noisy actions and the weights are entangled. This is typically addressed by learning an auxiliary model, referred to as the guidance network cep .
In this work, we propose a fundamentally different approach to diffusion-model-based offline RL. In addition to sampling the actions from the behavior policy $\mu(\mathbf{a}|s)$, we also sample the weights $w(s,\mathbf{a})$ using the same DM. As shown in Sec. 3, this enables a sampling procedure that effectively generates actions from the target policy $\pi(\mathbf{a}|s)$. In contrast to prior approaches that rely on training two distinct models—a DM and an external guidance network—our method removes the need of an external guidance network by embedding the guidance information within the diffusion model itself. We refer to this central contribution as Self-Weighted Guidance (SWG), detailed in Sec.3. Our contributions are summarized as follows
  • The derivation of the SWG method for exact guidance (Prop. 3.3 and Thm. 3.5 in Sec. 3).
  • An implementation setup of weights models tailored for SWG (Sec. 4).
  • A proof-of-concept for SWG’s exact guidance capabilities in toy experiments (Sec. 5.1).
  • Experimental validation for SWG’s competitive performance when compared to the state of the art in D4RL’s challenging environments, with a simpler training setup (Sec. 5.2).
  • Ablation studies exploring SWG’s robustness and scaling properties (Secs. 5.3 and 5.4).
2 Background
2.1 Offline reinforcement learning
Consider an agent that, following an unknown behavior policy $\mu(\mathbf{a}|s)$, interacted with the environment generating samples $(s,\mathbf{a},r,s^{\prime})$ recorded in a dataset $\mathcal{D}_{\mu}$. The goal of offline RL is to estimate the optimal policy $\pi(\mathbf{a}|s)$ from $\mathcal{D}_{\mu}$. Modeling the target policy as a weighted behavior policy in Eq. (1) is commonplace across various offline RL methods, some of which are presented next.
Constrained policy optimization peters_2010_constrained_policy ; AWRPeng19 ; nair2020awac regularizes the learned policy to ensure it remains close to the behavior policy by:
$\pi^{*}(\mathbf{a}|s)=\arg\max_{\pi}\mathbb{E}_{s\sim D_{\mu}}\left[\mathbb{E}_{\mathbf{a}\sim\pi(\mathbf{a}|s)}Q(s,\mathbf{a})-\frac{1}{\beta}D_{KL}(\pi(\cdot|s)\|\mu(\cdot|s))\right],$ (2)
where $Q(s,\mathbf{a})$ is the state-action function that indicates the quality of an action given a state, in terms of the expected future discounted rewards under a specific policy.
Following peters_2010_constrained_policy ; AWRPeng19 ; nair2020awac , it can be shown that the optimal policy $\pi^{*}(\mathbf{a}|s)$ in Eq. (2) has a weighted structure as in Eq. (1), given by:
$\pi^{*}(\mathbf{a}|s)\propto\mu(\mathbf{a}|s)e^{\beta Q{(s,\mathbf{a})}},$ (3)
where $Q(s,\mathbf{a})$ represents the Q-values function of the optimal policy. Notice that estimating the target policy Q-values requires querying out-of-sample actions.
Implicit Q-learning (IQL) iql avoids querying out-of-sample actions using an in-sample loss based on expectile regression. This way, it defines an implicit policy that outperforms the behavior policy. Furthermore, IQL employs advantage-weighted regression weighted_regression ; wang_weighted_regression ; AWRPeng19 ; nair2020awac to extract the policy, which can also be interpreted as a weighted behavior policy, given by:
$\pi(\mathbf{a}|s)\propto\mu(\mathbf{a}|s)e^{\beta A{(s,\mathbf{a})}},$ (4)
where $A{(s,\mathbf{a})}=Q(s,\mathbf{a})-V(s)$ is the advantage function learned with expectile regression.
Implicit Diffusion Q-learning (IDQL) idql generalizes implicit formulations to the use of an arbitrarily convex function $f$:
$V^{*}(s)=\arg\min_{V(s)}\mathbb{E}_{\mathbf{a}\sim\mu(\mathbf{a}|s)}\left[f(Q(s,\mathbf{a})-V(s))\right],$ (5)
thus defining a weighted behavior implicit policy given by
$\pi_{\text{imp}}(\mathbf{a}|s)\propto\mu(\mathbf{a}|s)w(s,\mathbf{a}),$ (6)
where $w(s,\mathbf{a})=\frac{|f^{\prime}(Q(s,\mathbf{a})-V^{*}(s))|}{|Q(s,\mathbf{a})-V^{*}(s)|}$. Furthermore, IDQL explores different convex functions $f$, such as expectiles iql , quantiles quantilekoenker and exponential functions.
2.2 Diffusion models in offline RL
Diffusion models (DM) sohl2015deep ; ddpm ; songscore approximate an unknown data distribution [^1] We denote data with the symbol $\mathbf{a}$ (instead of the standard $x$) for consistency, since in our setting the data are the agent’s actions. $q_{0}(\mathbf{a}_{0})$ using a dataset $\mathcal{D}$ by defining a sequence of $k\in\{0,\ldots,K\},K\in\mathbb{N},$ noisy distributions defined as follows:
$q_{k}(\mathbf{a}_{k}|\mathbf{a}_{0})\sim\mathcal{N}(\mathbf{a}_{k};\alpha_{k}\mathbf{a}_{0},\sigma_{k}^{2}\textbf{I}),$ (7)
where $\alpha_{k}$ and $\sigma_{k}$ are predefined noise schedules, s.t., $\mathbf{a}_{K}\sim\mathcal{N}(0,\mathbf{I})$ as $K\to\infty$. Therefore, samples from $q_{0}$ can be obtained by first sampling from $\mathcal{N}(0,\mathbf{I})$, and then solving the reverse diffusion process. In practice, the reverse diffusion is approximated by a model $\epsilon_{\theta}$ that predicts the noise to be removed from a noisy sample $\mathbf{a}_{K}\sim q_{K}$, to produce a target sample $\mathbf{a}_{0}\sim q_{0}$. The training loss for $\epsilon_{\theta}$ is ddpm
$\arg\min_{\theta}\mathbb{E}_{\mathbf{a}_{0},\epsilon_{0},k}||{\epsilon_{0}}-{\epsilon}_{{\theta}}(\mathbf{a}_{k},k)||_{2}^{2},$ (8)
where $\mathbf{a}_{0}\sim q_{0}$, $\epsilon_{0}\sim\mathcal{N}(0,\textbf{I})$, $k\sim\mathcal{U}(1,K)$ and $\mathbf{a}_{k}=\alpha_{k}\mathbf{a}_{0}+\sigma_{k}\epsilon_{0}$. The noise is related to the score function songscore of the intermediate noisy distribution as follows:
$\frac{-\epsilon_{0}}{\sigma_{k}}=\underbrace{\nabla_{\mathbf{a}_{k}}\log{q_{k}(\mathbf{a}_{k}})}_{\text{Score function}}.$ (9)
The weighted behavior policy expression in Eq. (1) admits using DMs in offline RL. To condition the DM on states, the noise prediction network can be made an explicit function of the state, that is, $\epsilon(\mathbf{a}_{k},k)=\epsilon(\mathbf{a}_{k},k,s)$. Previous approaches janner2022diffuser ; sfbc ; idql ; cep ; diff_dice fit a DM $\mu_{\theta}(\mathbf{a}|s)$ to the behavior policy $\mu(\mathbf{a}|s)$, and learn the weights $w(s,\mathbf{a})$ using methods described in Sec. 2.1. Then, samples from the target policy are obtained via resampling or guidance, explained as follows.
Resampling involves first sampling a batch of actions from the learned behavior policy $\mu_{\theta}(\mathbf{a}|s)$, and then resampling them using the weights as probabilities of a categorical distribution idql ; sfbc , in a way inspired by importance sampling. Alternatively, one can adopt a greedy policy that simply selects the action that maximizes the weight, with the aim of improving performance.
Guidance aims to obtain samples from the target policy by modifying the diffusion, adding a score term approximated by an external guidance network $g_{\phi}$ janner2022diffuser ; cep ; diff_dice . The $k$-step score in a diffusion sampling procedure is given by:
$\underbrace{\nabla_{\mathbf{a}_{k}}\log\pi_{k}(\mathbf{a}_{k}|s)}_{\text{target score}}\approx\underbrace{\nabla_{\mathbf{a}_{k}}\log\mu_{\theta}(\mathbf{a}_{k},k,s)}_{\approx\frac{-\epsilon_{\theta}(\mathbf{a}_{k},k,s)}{\sigma_{k}}}+\underbrace{\nabla_{\mathbf{a}_{k}}g_{\phi}(\mathbf{a}_{k},k,s)}_{\text{guidance score}},$ (10)
where $\{\mathbf{a}_{k}\}_{0}^{K}$ are the noisy versions of the actions, with $\mathbf{a}_{0}$ being the true noiseless action. As shown by cep , an unbiased estimate of the guidance score that effectively produces samples from the target distribution should satisfy:
$\nabla_{\mathbf{a}_{k}}g_{\phi}(\mathbf{a}_{k},k,s)=\nabla_{\mathbf{a}_{k}}\log\mathbb{E}_{\mu(\mathbf{a}_{0}|\mathbf{a}_{k},s)}[w(\mathbf{a}_{0},s)],$ (11)
where the right-hand term is intractable for most policies and weight functions. Guidance methods that are unbiased towards this objective are referred to as exact guidance methods cep .
Diffuser janner2022diffuser , trains the guidance network with a mean-squared-error objective that produces a biased target policy, in the sense that it does not match the form in Eq. (11). Conversely, exact guidance methods such as Contrastive Energy Prediction (CEP) cep use a contrastive learning objective that requires querying a batch of out-of-sample actions for each state in the dataset. Lastly, D-DICE diff_dice , proposes In-sample Guidance Learning (IGL) as an in-sample loss to learn the guidance network and uses a guide-then-select approach to obtain samples.
Remark 2.1.
While exact guidance methods effectively produce samples from the target distribution, they require learning an external network $g_{\phi}$ through an intricate loss function $\mathcal{L}_{g}(\phi)$. This motivates our proposal called Self-Weighted Guidance, which leverages DM’s without resorting to a special training objective to conduct exact guidance.
3 Self-weighted guidance
3.1 Problem statement
Let us consider a dataset $\mathcal{D}_{A}=\{\mathbf{a}^{(i)}\}_{i=1}^{N}\subset\mathbb{R}^{d}$ of i.i.d. realizations of a random variable $A\sim q(\mathbf{a})$; these will be the agent’s actions. Recall that we aim to sample from the unnormalized target $p(\cdot)$ that relates to $q(\cdot)$ as
$p(\mathbf{a})\propto q(\mathbf{a})w(\mathbf{a}),$ (12)
where $w:\mathbb{R}^{d}\to\mathbb{R}$ is a known [^2] Though in offline RL the weight function $w$ has to be learned, we consider $w$ known for now. weight function. Sampling can be achieved by fitting a DM, denoted $\epsilon_{\theta}$, to $q(\mathbf{a})$ trained on $\mathcal{D}_{A}$. Then, in the $k$-th denoising step, the score of the target distribution (referred to as the target score) satisfies:
$\underbrace{\nabla_{\mathbf{a}_{k}}\log p_{k}(\mathbf{a}_{k})}_{\text{target score}}=\underbrace{\nabla_{\mathbf{a}_{k}}\log q_{k}(\mathbf{a}_{k})}_{\approx\frac{-\epsilon_{\theta}(\mathbf{a}_{k},k)}{\sigma_{k}}}+\underbrace{\nabla_{\mathbf{a}_{k}}\log{\mathbb{E}_{q(\mathbf{a}_{0}|\mathbf{a}_{k})}\left[w(\mathbf{a}_{0})\right]}}_{\text{intractable score}},$ (13)
where $\mathbf{a}_{0}\in\mathcal{D}_{A}$ is a data sample, $\mathbf{a}_{1},\ldots,\mathbf{a}_{K}$ are the noisy copies of the data on which the DM is trained, and the subindex $\cdot_{k}$ in the distributions denote the denoising step. Though $\nabla_{\mathbf{a}_{k}}\log q_{k}(\mathbf{a}_{k})$, referred to as intermediate score, is directly provided by the DM, estimating the target score in Eq. (13) is challenging due to the intractability of the right-hand term in the general case cep .
3.2 Proposal: diffusing over actions and weights
Let us consider a random variable (RV) $Z=(A,W)$ given by the concatenation of $A$, describing the actions in the previous section, and $W$, which is given by passing $A$ through the weight function $w(\cdot)$. Since $w(\cdot)$ is a deterministic function of $A$, the distribution of $Z$ is given by the pushforward measure of the distribution of $A$ through the mapping $T:\mathbf{a}\mapsto(\mathbf{a},w(\mathbf{a}))$. With a slight abuse of notation, we denote the pushforward measures corresponding to $p$ and $q$ in Eq. (12) with the same symbols, i.e., $p(\mathbf{z})=T_{\#}p(\mathbf{a})$ and $q(\mathbf{z})=T_{\#}q(\mathbf{a})$, respectively. Therefore, we have
$p(\mathbf{z})$ $=p(\mathbf{a},w)=p(w|\mathbf{a})p(\mathbf{a})=p(\mathbf{a})\delta_{w=w(\mathbf{a})}$ (14)
$q(\mathbf{z})$ $=q(\mathbf{a},w)=q(w|\mathbf{a})q(\mathbf{a})=q(\mathbf{a})\delta_{w=w(\mathbf{a})},$ (15)
which, combined with Eq. (12), give $p(\mathbf{z})\propto q(\mathbf{z})w$. In this new formulation, $w$ is no longer to be considered as a function of a, but rather as a variable for which there are available observations. For clarity of presentation, we define the following function
Definition 3.1 (Extraction function).
The function $\phi_{w}:\mathbb{R}^{d+1}\to\mathbb{R}$ extracts the weight component from the concatenated variable $\mathbf{z}=[\mathbf{a},w]$. That is, $\phi_{w}(\cdot)=\phi_{w}([\mathbf{a},w])=w$.
Remark 3.2.
The extraction function allows us to express
$p(\mathbf{z})\propto q(\mathbf{z})\phi_{w}(\mathbf{z}),$ (16)
thus implying that the target distribution for $\mathbf{z}$ follows the same structure as the original distribution in Eq. (12). This motivates jointly learning weights and actions with the same DM. Since the extraction function $\phi_{w}(\mathbf{z})$ is linear, it is feasible to estimate the (otherwise intractable) score within our DM.
We can now implement a DM to sample from the augmented variable $Z$. To this end, we first extend the dataset $\mathcal{D}_{A}=\{(\mathbf{a}^{(i)})\}_{i=1}^{N}$ to $\mathcal{D}_{Z}=\{(\mathbf{a}^{(i)},w^{(i)})\}_{i=1}^{N}\subset\mathbb{R}^{d+1}$, where $w^{(i)}=w(\mathbf{a}^{(i)})$ is a weight observation. Notice that the dataset $\mathcal{D}_{Z}$ provides access to the weight function only on the support of the original data in $\mathcal{D}_{A}$. For consistency, we also denote realizations of $Z$ as $\mathbf{z}^{(i)}=(\mathbf{a}^{(i)},w^{(i)})$.
Therefore, we define a $K$-step diffusion with noise schedule $\{\alpha_{k}\}_{k=1}^{K}$, $\{\sigma_{k}\}_{k=1}^{K}$, and train a noise-prediction DM ${\epsilon}_{\theta}$ for $Z$ using the augmented dataset $\mathcal{D}_{Z}$, minimizing the standard diffusion loss in Eq. (8). In our notation, this is
$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{z}_{0},\epsilon_{0},k}||{\epsilon_{0}}-{\epsilon}_{{\theta}}(\mathbf{z}_{k},k)||_{2}^{2},$ (17)
where $\mathbf{z}_{0}\sim\mathcal{D}_{Z}$, $\mathbf{z}_{k}=\alpha_{k}\mathbf{z}_{0}+\sigma_{k}\epsilon_{0}$, $\epsilon_{0}\sim\mathcal{N}(0,\textbf{I})$, and $k\sim\mathcal{U}(1,K)$. Algorithm 1 describes our proposed DM training procedure.
Algorithm 1 Training of joint DM over weights and actions
Input: Dataset $\mathcal{D}_{A}$, weight function $w(\cdot)$, number of diffusion steps $K$, noise schedule $\{\alpha_{k}\}_{k=1}^{K}$, $\{\sigma_{k}\}_{k=1}^{K}$, learning rate $\lambda$, and batch size $B$.
Initialize: $\theta$ parameters for the diffusion model ${\epsilon}_{\theta}$
Augment dataset $\mathcal{D}_{A}\to\mathcal{D}_{Z}$ as described in Sec. 3.2.
for each parameters update do
     Draw $B$ samples $\mathbf{z}_{0}\sim\mathcal{D}_{Z}$ and $B$ indices $k\sim\mathcal{U}(1,K)$
     Add noise to each sampled data $\mathbf{z}_{0}$ for respective $k$ time as $\mathbf{z}_{k}=\alpha_{k}\mathbf{z}_{0}+\sigma_{k}\epsilon_{0}$, $\epsilon_{0}\sim\mathcal{N}(0,\textbf{I})$
     Calculate loss $L(\theta)\approx\frac{1}{B}\sum_{B}||{\epsilon_{0}}-{\epsilon}_{{\theta}}(\mathbf{z}_{k},k)||_{2}^{2}$
     Update parameters $\theta\leftarrow\theta-\lambda\nabla_{\theta}L(\theta)$
end for
Return: Trained diffusion model ${\epsilon}_{\theta}$
3.3 Augmented exact guidance
Proposition 3.3.
The required score to sample from the target distribution in Eq. (16) is given by
$\nabla_{\mathbf{z}_{k}}\log p_{k}(\mathbf{z}_{k})=\underbrace{\nabla_{\mathbf{z}_{k}}\log q_{k}(\mathbf{z}_{k})}_{\approx\frac{-\epsilon_{\theta}(\mathbf{z}_{k},k)}{\sigma_{k}}}+\underbrace{\nabla_{\mathbf{z}_{k}}\log{\mathbb{E}_{q(\mathbf{z}_{0}|\mathbf{z}_{k})}\left[\phi_{w}(\mathbf{z}_{0})\right]}}_{\text{intractable term}}.$ (18)
The proof for Proposition 3.3 follows as an extension of cep to our setting (see Appendix A.1).
Remark 3.4.
Since the extended formulation in Eq. (16) is defined over $\mathbf{z}$, the gradients required to implement guidance in Eq. (18) are taken with respect to $\mathbf{z}_{k}$, rather than $\mathbf{a}_{k}$ (as in Eq. (13)). This is a desired feature, since the actions and weights are highly coupled and thus guiding them separately can lead to inefficient sampling.
The key insight of our proposal is that the intractable term in Eq. (18) can be recovered directly from the DM described above using the so-called data prediction formula efron2011tweedie ; variational_diffusion_models given by:
$\mathbb{E}_{q(\mathbf{z}_{0}|\mathbf{z}_{k})}[\mathbf{z}_{0}]\approx\mathbf{z}_{\theta}(\mathbf{z}_{k},k):=\frac{\mathbf{z}_{k}-\sigma_{k}\epsilon_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}.$ (19)
Using Eq. (19), and since the extraction function $\phi_{w}(\cdot)$ is linear, the target score $\nabla_{\mathbf{z}_{k}}\log p_{k}(\mathbf{z}_{k})$ in Eq. (18) can be computed with the proposed DM, as stated in Theorem 3.5 (see the proof in Appendix A.2).
Theorem 3.5.
Let $\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)$ be the DM found as the global minimizer of Eq. (17). The exact target score of Eq. (18) is given by:
$\nabla_{\mathbf{z}_{k}}\log p_{k}(\mathbf{z}_{k})=\underbrace{\nabla_{\mathbf{z}_{k}}\log q_{k}(\mathbf{z}_{k})}_{\frac{-\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\sigma_{k}}}+\underbrace{\nabla_{\mathbf{z}_{k}}\log{\left(\phi_{w}\left(\frac{\mathbf{z}_{k}-\sigma_{k}\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}\right)\right)}}_{\text{self guidance}},$ (20)
where recall that $\phi_{w}(\mathbf{z})=w$ simply extracts the weight component from $\mathbf{z}=[\mathbf{a},w]$.
Our method is referred to as Self-Weighted Guidance (SWG), since the DM samples from the target distribution $\mathbf{z}_{0}\sim p(\mathbf{z}_{0})$ by guiding itself, that is, the guidance signal is a function of the same model (see Algorithm 2). SWG provides an elegant way of performing exact guidance, since it computes gradients directly on the DM $\epsilon_{\theta}$ instead of an external guidance network with a dedicated —and possibly intricate— loss function (see Appendix C). This translates to not having to train a separate guidance network, thereby simplifying the training pipeline. Table 1 presents a conceptual comparison between SWG and other DM-based guidance approaches to offline RL.
Table 1: SWG in context: differences with other DM-based guidance approaches to offline RL.
Exact guidance Requires External Guidance Network
Diffuser janner2022diffuser No Yes
CEP cep Yes Yes
D-DICE diff_dice Yes Yes
SWG (ours) Yes No
Algorithm 2 Self-Weighted Guidance sampling
Input: Pretrained DM ${\epsilon}_{\theta}$, number of diffusion steps $K$, and noise schedule $\{\alpha_{k}\}_{k=1}^{K}$, $\{\sigma_{k}\}_{k=1}^{K}$.
Sample noise $\mathbf{z}_{k}\sim\mathcal{N}(0,\textbf{I})$
for each $k$ in reversed $0:K$ do
     Forward pass on diffusion model ${\epsilon}_{\theta}(\mathbf{z}_{k},k)$
     $\hat{\epsilon}:={\epsilon}_{\theta}(\mathbf{z}_{k},k)-\sigma_{k}\nabla_{\mathbf{z}_{k}}\log{(\phi_{w}(\frac{\mathbf{z}_{k}-\sigma_{k}{\epsilon}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}))}$
     Calculate $\mathbf{z}_{k-1}$ with noise prediction $\hat{\epsilon}$
     Set $\mathbf{z}_{k}:=\mathbf{z}_{k-1}$
end for
Return: Sample $\mathbf{z}_{0}$
4 Implementing SWG in offline RL
There are two main practical aspects regarding the implementation of SWG in the offline RL setup. First, recall that for notational and conceptual simplicity, SWG was presented in a way independent of the states. In practice, however, state-dependent DM’s are trained by considering the state as an input to the noise predictor network, that is, $\epsilon(\mathbf{z}_{k},k)=\epsilon(\mathbf{z}_{k},k,s)$ in Eq. (17). Second, SWG was introduced in Sec. 3 assuming a known weight function $w$, however, in offline RL the weights need to be estimated beforehand. We next present how to adapt and implement different weight formulations within the proposed SWG.
Expectile. The IDQL idql $\tau$-expectile weight is defined as $w_{\tau}^{2}(s,\mathbf{a})=\left|\tau-1_{\{Q(s,\mathbf{a})<V_{\tau}^{2}(s)\}}\right|$, where $V_{\tau}^{2}(s)$ is the value function learned with the expectile loss, and $1_{\{\cdot\}}$ is the indicator function. Since this weight takes only two possible values, $w_{\tau}^{2}(s,\mathbf{a})\in\{\tau,1-\tau\}\subset[0,1]$, it may fail to distinguish between similar actions by assigning them identical weights. To address this limitation, we propose a smooth variant of these weights, given by
$w(s,\mathbf{a})=\left(\frac{1}{1+e^{-A(s,\mathbf{a})}}\right)\tau+\left(1-\frac{1}{1+e^{-A(s,\mathbf{a})}}\right)(1-\tau).$ (21)
Where we apply a sigmoid function to the advantage $A(s,\mathbf{a})$ to obtain a continuous weight function. This weight function is utilized in our D4RL experiments (Sec.5.2), and the full training procedure for the value functions is detailed in Algorithm 3 in Appendix D.2.
Exponential. From the same value functions learned with expectile loss, we can define an exponential weight model derived from the IQL iql policy extraction method as $w(s,\mathbf{a})=e^{\beta A(s,\mathbf{a})}$, where $A(s,\mathbf{a})=Q(s,\mathbf{a})-V(s)$ is the advantage function. Though this formulation addresses the sparsity issue arising in the expectile weight, its exponential form could lead to arbitrarily large values, and is thus potentially unstable during training of SWG. Using the advantage function can help mitigate instabilities, as the value function serves as a normalizer for the Q-values.
Linex. Another exponential formulation can be obtained when learning the value function using the linex function $lx:u\mapsto\exp(u)-u$. As found by idql , this defines the following weight:
$w_{\text{exp}}(s,\mathbf{a})=\frac{\alpha\big{|}\exp\big{(}\alpha\big{(}Q(s,\mathbf{a})-V_{\text{exp}}(s)\big{)}\big{)}-1\big{|}}{\big{|}Q(s,\mathbf{a})-V_{\text{exp}}(s)\big{|}},$ (22)
where $V_{\text{exp}}(s)$ is the value function learned with the linex loss, and $\alpha$ is a temperature coefficient.
We emphasize that, among the various weighted behavior policies in the literature, we focus on avoiding querying out-of-sample actions, as they can lead to overestimation errors iql . Sec. 5.3 provides an experimental comparison of different weight formulations within SWG on D4RL, with further details in Appendix D.2.
Remark 4.1.
The standard practice across most offline RL methods is to estimate $Q(s,\mathbf{a})$ and $V(s)$ to then construct the weight function; this adds complexity to the overall training pipeline. We clarify that though SWG does not require an external guidance network during sampling, it still requires learning of the weight function prior to sampling.
Lastly, in our implementation the strength of the guidance score is controlled via a hyperparameter $\rho$, called guidance scale diffusion_gans , as follows $\hat{\epsilon}:={\epsilon}_{\theta}(\mathbf{z}_{k},k)-\rho\sigma_{k}\nabla_{\mathbf{z}_{k}}\log{(\phi_{w}(\frac{\mathbf{z}_{k}-\sigma_{k}{\epsilon}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}))}$.
5 Experiments
The proposed SWG was implemented and validated through a set of experiments targeting the following questions (relevant experiments for each question in parentheses):
  1. 1.
    Can SWG sample from the target distribution? (5.1)
  2. 2.
    How does SWG perform against the state-of-the-art in the D4RL benchmark? (5.2)
  3. 3.
    How do different weight formulations affect SWG? (5.3)
  4. 4.
    How does SWG’s inference time scale with increasing model size? (5.4)
The experimental setup and benchmarks are in Appendix D, additional hyperparameter sensitivity experiments are in Appendices D.3 and E, and our code can be found in SWG repository.
5.1 Sampling from the target distribution
We considered the toy experiments from cep , which use a weight function $w(\mathbf{a})=\exp(-\beta\xi(\mathbf{a}))$, where $\xi(\mathbf{a})$ is an energy function, and $\beta$ is the inverse temperature coefficient. Figure 1 shows SWG samples from a spiral distribution and offers a qualitative assessment. The samples closely matching the ground truth for different values of $\beta$ suggest that SWG was able to sample from the target distribution, thus learning the relationship between the data distribution and the weights. The reader is referred to Appendix E for more toy examples of this kind.
Figure 1: SWG applied to a spiral distribution for different temperature coefficients $\beta$. Top: samples generated by SWG. Bottom: ground truth samples. When $\beta=0$, we have $w(\mathbf{a})=1$ and thus the target distribution coincides with the data distribution.
5.2 D4RL experiments
Using the D4RL benchmark fu2020d4rl , we compared SWG to classical offline RL methods (CQL cql and IQL iql ), diffusion-based resampling methods (SfBC sfbc and IDQL idql ), diffusion-guided methods (Diffuser janner2022diffuser , D-DICE (IGL) diff_dice , and GQPO (CEP)cep ), and to D-QL dql , which trains a DM using Q-values as regularization term.
Our experiments are performed across four task groups: 1) Locomotion is the main offline RL benchmark and primarily involves moving an agent across a plane. 2) Ant Maze is considered a more challenging task, where the goal is to guide an ant agent through a maze to reach a target. 3) Franka Kitchen involves goal-conditioned manipulation in a realistic kitchen environment. Lastly, 4) Adroit Pen focuses on high-dimensional dexterous manipulation to reorient a pen.
These experiments were conducted with the expectile weight proposed in IDQL idql modified as detailed in Sec. 4, and using $K=15$ diffusion steps. See Appendix D.2 for more details.
The results shown in Table 2 reveal that SWG performance was on par with state-of-the-art DM’s in challenging environments such as Ant Maze, Franka Kitchen and Adroit Pen tasks. In locomotion tasks, however, SWG achieved average performance with respect to previous methods. Nevertheless, the locomotion benchmark is considered to be saturated rafailovd5rl and does not necessarily represent a trustworthy criterion. Also note that SWG outperformed classical offline RL methods such as CQL and IQL in most tasks.
Based on the D4RL results, we argue that SWG offers a tradeoff between having a simpler training setup with fewer components than current DM guidance-based approaches at the cost of a lower performance with respect to the top-performing method, D-DICE diff_dice .
Table 2: SWG against the state-of-art on the D4RL benchmark: Mean and standard deviation of the normalized returns fu2020d4rl under $5$ random seeds. Values within $5\%$ of the best performing method in each task (row) are in bold font. See Appendix D for full experimental details.
Dataset Environment CQL IQL SfBC IDQL Diffuser D-QL D-DICE QGPO SWG (ours)
Medium-Expert HalfCheetah $62.4$ $86.7$ $\bf{92.6}$ $\bf{95.9}$ $79.8$ $\bf{96.8}$ $\bf{97.3}$ $\bf{93.5}$ $\bf{92.8}\pm 0.6$
Medium-Expert Hopper $98.7$ $91.5$ $\bf{108.6}$ $\bf{108.6}$ $\bf{107.2}$ $\bf{111.1}$ $\bf{112.2}$ $\bf{108.0}$ $\bf{109.9}\pm 3.2$
Medium-Expert Walker2d $\bf{111.0}$ $\bf{109.6}$ $\bf{109.8}$ $\bf{112.7}$ $\bf{108.4}$ $\bf{110.1}$ $\bf{114.1}$ $\bf{110.7}$ $\bf{110.7}\pm 1.7$
Medium HalfCheetah $44.4$ $47.4$ $45.9$ $51.0$ $44.2$ $51.1$ $\bf{60.0}$ $54.1$ $46.9\pm 0.2$
Medium Hopper $58.0$ $66.3$ $57.1$ $65.4$ $58.5$ $90.5$ $\bf{100.2}$ $\bf{98.0}$ $68.6\pm 4.4$
Medium Walker2d $79.2$ $78.3$ $77.9$ $82.5$ $79.7$ $\bf{87.0}$ $\bf{89.3}$ $\bf{86.0}$ $74.0\pm 2.5$
Medium-Replay HalfCheetah $46.2$ $44.2$ $37.1$ $45.9$ $42.2$ $\bf{47.8}$ $\bf{49.2}$ $\bf{47.6}$ $42.5\pm 1.5$
Medium-Replay Hopper $48.6$ $94.7$ $86.2$ $92.1$ $96.8$ $\bf{101.3}$ $\bf{102.3}$ $96.9$ $73.9\pm 2.3$
Medium-Replay Walker2d $26.7$ $73.9$ $65.1$ $85.1$ $61.2$ $\bf{95.5}$ $\bf{90.8}$ $84.4$ $60.6\pm 4.1$
Average (Locomotion) $63.9$ $76.9$ $75.6$ $82.1$ $75.3$ $\bf{88.0}$ $\bf{90.6}$ $\bf{86.6}$ $75.5$
Default AntMaze-umaze $74.0$ $87.5$ $92.0$ 94.0 - $\bf{93.4}$ $\bf{98.1}$ $\bf{96.4}$ $92.4\pm 2.6$
Diverse AntMaze-umaze $\bf{84.0}$ $62.2$ $\bf{85.3}$ 80.2 - $66.2$ $\bf{82.0}$ $74.4$ ${70}\pm 4.2$
Play AntMaze-medium $61.2$ $71.2$ $81.3$ 84.5 - $76.6$ $\bf{91.3}$ $83.6$ $\bf{89.6}\pm 2.3$
Diverse AntMaze-medium $53.7$ $70.0$ $82.0$ 84.8 - $78.6$ $\bf{85.7}$ $\bf{83.8}$ $\bf{88.2}\pm 5.2$
Play AntMaze-large $15.8$ $39.6$ $59.3$ 63.5 - $46.4$ $\bf{68.6}$ $\bf{66.6}$ $55.4\pm 4.6$
Diverse AntMaze-large $14.9$ $47.5$ $45.5$ 67.9 - $56.6$ $\bf{72.0}$ $64.8$ $55.2\pm 3$
Average (Ant Maze) $50.6$ $63.0$ $74.2$ 79.1 - $69.6$ $\bf{83}$ $78.3$ $75.1$
Human Adroit Pen $37.5$ $71.5$ $76$ $72.8$ $\bf{84.4}$ $72.8$ $78.5\pm 6.3$
Cloned Adroit Pen $39.2$ $37.3$ $64$ $57.3$ $\bf{83.8}$ $54.2$ $62.3\pm 6.7$
Average (Adroit Pen) $38.4$ $54.4$ $70$ $65.1$ $\bf{84.1}$ $63.5$ $70.4$
Partial Franka Kitchen $49.8$ $46.3$ $47.9$ $60.5$ $\bf{78.3}$ $69.0\pm 4.9$
Mixed Franka Kitchen $51.0$ $51.0$ $45.4$ $62.6$ $\bf{67.8}$ $61.0\pm 2.2$
Complete Franka Kitchen $43.8$ $62.5$ $77.9$ $\bf{84.0}$ $72.3\pm 2.6$
Average (Franka Kitchen) $48.2$ $53.3$ $57.1$ $\bf{69.0}$ $\bf{67.4}$
5.3 Ablation study for different weights
We then evaluated SWG’s performance under different weights formulations. We considered the Linex, Expectile and the exponential expectile weights presented in Sec. 4. This test was conducted on D4RL locomotion, using the same model and hyperparameter tuning to ensure a fair comparison, and the results are shown in Table 3. The expectile weight yielded the best performance, the exponential expectile showed moderate performance, and the Linex weight performed substantially worse. We conjecture that this is due to training instability introduced by the exponential weights in general. Most importantly, these results are consistent with the findings reported in IDQL idql .
Table 3: Ablation study for different weights models implemented alongside SWG on the locomotion D4RL benchmark.
Expectile Exp-expectile Linex
Average (locomotion) $\bf{75.5}$ $69.9$ $63.6$
5.4 How does SWG’s inference time scale with its size?
Since SWG requires computing gradients through the DM, we investigated its scalability properties via two experiments examining both depth and width scaling of its networks. For depth scaling, we used a simple multi-layer perceptron (MLP) with a hidden dimension of 2056 and measured the inference time across varying numbers of layers. For width scaling, we used a fixed 4-layer MLP with varying hidden dimension, measuring average inference time accordingly (See Appendix D.5).
As shown in Figures 3 and 3, SWG exhibited a linear scaling with respect to model depth and exponential scaling with respect to the width scaling (hidden dimension). Additionally, and as expected, sampling with guidance (i.e., computing gradients) translates to larger inference times. However, qualitatively, we can observe that both sampling methods scale at a similar rate. Note that scaling issues related to gradient computation can be easily avoided by designing the DM to output the weight component in its early layers. This allows gradients to be computed with respect to a fixed module, without requiring backpropagation through the entire model.
Figure 2: SWG depth scaling: inference time (s) vs number of layers (mean $\pm$ 2 standard deviations over 10,000 runs).
Figure 3: SWG width scaling: inference time (s) vs hidden dimension (mean $\pm$ 2 standard deviations over 10,000 runs).
6 Conclusion
We have proposed a joint DM over actions and weights for optimal policy discovery in offline RL. Our approach addresses a shortcoming of current DM-based methods, which struggle to compute the guidance score network. In our proposal, the computation of the scores is straightforward, as they are readily available from the DM implementation. The proposed method, termed Self-Weighted Guidance (SWG), owes its name to having the weight guidance information embedded into the DM. Compared to previous methods, SWG conducts exact guidance using a streamlined setup which requires a single DM, and provides novel insights arising from modeling weights as random variables, thus demonstrating one more time the flexible modeling capabilities of DM’s.
Via toy examples and the standard D4RL benchmark, we have provided experimental validation of SWG’s ability to: i) effectively produce samples of the target distribution, and ii) perform comparably against state-of-the-art DMs on challenging environments such as Ant Maze, Franka Kitchen and Adroit Pen, while having a simpler training setup. Furthermore, we have provided an ablation study for different weight formulations within SWG, and also studied how SWG’s inference time scales with its network’s width and depth. We stress that the main advantage of SWG stems from its simplicity and reduced training setup compared with previous guidance methods.
Lastly, we claim that the proposed DM-based sampling methodology, which is guided by the same DM, has the potential for applications beyond the offline RL setting. Tasks involving weighted distributions such as controllable image and text generation diffusion_gans ; Li-2022-DiffusionLM can be addressed by the proposed SWG approach.
Limitations and future work.
Though SWG’s performed competitively on D4RL, it was outperformed by some recent diffusion methods released in the last two years, namely QGPO and D-DICE. We conjecture that this first version of SWG can be enhanced with tailored weight models and more sophisticated ODE solvers. Also, following our preliminary study on SWG’s scalability in Sec. 5.4 regarding gradient computation during inference, we claim that extracting the weight component from the first layers of the DM network will critically alleviate computational complexity.
Broader impacts.
This work presents a promising methodology toward offline RL, where policies can be learned from a controlled offline dataset. However, as with most works on diffusion models, potential misuse includes the generation of harmful data, such as fake images, videos, or audio.
Appendix A Proofs and derivations
A.1 Proof of Proposition 3.3: computation of the target score
We follow the derivation utilized in [31].
Let us express the normalizing constant of $p_{0}(\mathbf{z}_{0})$ as
[…]
Recall that $\phi_{w}(\cdot)$ extracts the $w$ component of $z$. Then, we have:
$p_{0}(\mathbf{z}_{0})=\frac{q_{0}(\mathbf{z}_{0})\phi_{w}(\mathbf{z}_{0})}{A}.$
According to the diffusion models definition, we have:
[…]
and finally
$\nabla_{\mathbf{z}_{k}}\log p_{k}(\mathbf{z}_{k})=\nabla_{\mathbf{z}_{k}}\log q_{k}(\mathbf{z}_{k})+\nabla_{\mathbf{z}_{k}}\log{\mathbb{E}_{q_{0}(\mathbf{z}_{0}|\mathbf{z}_{k})}\left[\phi_{w}(\mathbf{z}_{0})\right]}.$
A.2 Proof of Theorem 3.5
Recall that we want to sample from $p(\mathbf{z})$ in Eq. (18).
Let us assume that the weight function $w(\cdot)$ is known, so we have access to a labeled dataset consisting of samples $\mathbf{z}_{0}=(\mathbf{a}_{0},w(\mathbf{a}_{0}))$. We suppose a $K$ step diffusion process over $\mathbf{z}_{0}$, as $\mathbf{z}_{k}=\alpha_{k}\mathbf{z}_{0}+\sigma_{k}\epsilon_{0}$, $\epsilon\sim\mathcal{N}(0,\textbf{I})$.
To tackle the problem in Equation (10), let us assume we train a noise predicting diffusion model on the data $\mathbf{z}_{0}$, $\epsilon_{\theta}(\mathbf{z}_{k},k)$, $k\in[0,K]$, parametrized by $\theta$. Then, the optimal diffusion model $\epsilon^{*}$ is defined by the following optimization problem.
$\operatorname*{argmin}_{{\theta}}||{\epsilon}_{0}-{\epsilon}_{{\theta}}(\mathbf{z}_{k},k)||_{2}^{2}.$ (23)
Let us also recall the data prediction formula [8, 19] for diffusion models, which states that the optimal diffusion model satisfies:
$\mathbb{E}_{q_{0k}(\mathbf{z}_{0}|\mathbf{z}_{k})}[\mathbf{z}_{0}]\approx z^{*}_{\theta}(\mathbf{z}_{k},k):=\frac{\mathbf{z}_{k}-\sigma_{k}\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}.$ (24)
We can use this formula to compute the intractable score in Eq. (18). Since the extraction function $\phi_{w}(\cdot)$ is linear (see Def. 3.1 in Sec. 3), we can swap the expectation and $\phi_{w}$ to obtain
$\mathbb{E}_{q_{0k}(\mathbf{z}_{0}|\mathbf{z}_{k})}[\phi_{w}(\mathbf{z}_{0})]=\phi_{w}(\mathbb{E}_{q_{0k}(\mathbf{z}_{0}|\mathbf{z}_{k})}[\mathbf{z}_{0}])=\phi_{w}\left(\frac{\mathbf{z}_{k}-\sigma_{k}\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}\right).$ (25)
Finally, plugging this term in Eq. (18)
$\nabla_{\mathbf{z}_{k}}\log p_{k}(\mathbf{z}_{k})=\underbrace{\nabla_{\mathbf{z}_{k}}\log q_{k}(\mathbf{z}_{k})}_{\frac{-\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\sigma_{k}}}+\underbrace{\nabla_{\mathbf{z}_{k}}\log{\left(\phi_{w}\left(\frac{\mathbf{z}_{k}-\sigma_{k}\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}\right)\right)}}_{{\text{self guidance}}}.$ (26)
Note that if we utilize a data prediction diffusion model $\mathbb{E}_{q_{0k}(\mathbf{z}_{0}|\mathbf{z}_{k})}[\mathbf{z}_{0}]=z^{*}_{\theta}(\mathbf{z}_{k},k)$ the target score can be expressed as:
$\nabla_{\mathbf{z}_{k}}\log p_{k}(\mathbf{z}_{k})=\underbrace{\nabla_{\mathbf{z}_{k}}\log q_{k}(\mathbf{z}_{k})}_{\frac{-\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\sigma_{k}}}+\underbrace{\nabla_{\mathbf{z}_{k}}\log{(\phi_{w}(z^{*}_{\theta}(\mathbf{z}_{k},k)))}}_{{\text{self guidance}}}.$ (27)
Notice we must make the assumption that $\phi_{w}(\frac{\mathbf{z}_{k}-\sigma_{k}\epsilon^{*}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}})>0$, which the optimal diffusion model satisfies if the weight function is positive $w(\mathbf{a})>0$ $\forall\mathbf{a}\in supp(A)$.
Appendix B Related work
In this section, we review prior work in key areas relevant to our approach, as a complement to Sec. 2, to better contextualize our contributions.
B.1 Offline RL
Offline RL methods must address common challenges such as overestimation errors for out-of-sample actions and distributional shift. Distributional shift is tackled by constraining the learned policy to remain close to the behavior policy, typically by adding a regularization term [25, 10, 35]. This regularization term can be utilized to learn the policy directly [10, 55, 25] or through a weighted regression procedure [41, 39, 35, 54].
Recent studies have proposed various methods for handling overestimation errors. For example, Conservative Q-Learning (CQL) [26] penalizes Q values of out-of-distribution actions to prevent overestimation. Implicit Q-Learning (IQL) [24] uses expectile regression to learn value functions, completely avoiding out-of-sample actions during training.
B.2 Diffusion models
Diffusion Models [48, 14, 50] are a recent powerful class of generative models that excel across numerous tasks, including, but not limited to, image synthesis [44, 7], text-to-image generation [43, 46] and video generation [16]. Controllable sampling has emerged as a necessary technique to ensure that generated samples possess desired features. Methods like Classifier Guidance [7] and Classifier-Free Guidance [15] aim to generate samples from the conditional distribution. Diffusion Models are limited by their slow sampling speed, as they require an autoregressive forward pass on a model. Existing attempts to improve sampling speed include: changing the Markovian formulation of Diffusion Models [49], distilling the model [47], and utilizing fast ODE solvers [33, 32].
B.3 Generative models in offline RL
Generative models have shown great promise in the field of Offline Reinforcement Learning (RL). Diffuser [17] performs planning with a diffusion model and improves results with a guidance network. Similarly, [1] proposes a method that utilizes conditional sampling to enhance results, employing Classifier-Free Guidance [15] to achieve this.
Diffusion policy methods have also been explored extensively. Diffusion-QL [53] adds Q-values as a regularization term during diffusion training to improve policy performance. Other notable methods include Implicit Diffusion Q-Learning (IDQL) [12] and SfBC [4] that model the behavior policy and simply utilize resampling from a batch of action candidates to improve performance. CEP [31], and D-DICE [34] are proposed as exact guidance methods, utilizing a diffusion model and an external guidance network to obtain samples from the target policy.
In addition to diffusion models, other generative models have been successfully applied to offline RL. Transformers, such as Decision Transformer (DT) [5] and Trajectory Transformer (TT) [18], have demonstrated impressive performance in various offline RL tasks. Variational Autoencoders (VAEs) [21] have also been employed to model the policy [10, 25].
Appendix C Relationship to previous guidance methods
In this section, we provide an analysis of previous guidance methods alongside our own, highlighting theoretical differences and potential advantages.
C.1 Comparison to previous guidance methods in offline RL
Let us consider the two existing exact-guidance DM methods in the offline RL literature: CEP [31] and IGL (from D-DICE [34]). Assuming a known weight function $w(\cdot)$, CEP learns the guidance network $g_{\theta}$ using the loss
$\mathcal{L}_{CEP}(\theta)=\mathbb{E}_{k,\mathbf{a}_{0,k}^{(1:M)}}\Bigg{[}-\sum_{i=1}^{M}w(s,\mathbf{a}_{0}^{(i)})\log\tfrac{e^{g^{(i)}_{\theta,k}}}{\sum_{j=1}^{M}e^{g^{(j)}_{\theta,k}}}\Bigg{]},$ (28)
where $k\sim\mathcal{U}(1,K)$, and for ease of notation $g^{(i)}_{\theta,k}=g_{\theta}(s,\mathbf{a}_{k}^{(i)},k)$, and $\mathbf{a}_{0,k}^{(1:M)}=\{\mathbf{a}_{j}^{(i)},\text{s.t.},j\in\{0,k\},i\in\{1,\ldots,M\}\}$. Recall that $\mathbf{a}_{0}$ is the data and $\mathbf{a}_{k}$ are the noisy copies. Since Eq. (28) requires $M>1$ action samples from each state in the dataset, its evaluation is likely to need out-of-sample actions, thus leading to overestimation errors [24, 34]. CEP [31] address this by pretraining a DM on the behavior policy and generating a fake action dataset containing $M>1$ samples for each state in the dataset.
On the other hand, IGL does not require querying out-of-sample actions using the following in-sample loss:
$\mathcal{L}_{IGL}(\theta)=\mathbb{E}_{k,\mathbf{a}_{0,k}}\Bigg{[}w(s,\mathbf{a}_{0})e^{-g_{\theta}(s,\mathbf{a}_{k},k)}+g_{\theta}(s,\mathbf{a}_{k},k)\Bigg{]},$ (29)
where $k\sim\mathcal{U}(1,K)$, $\mathbf{a}_{0}\sim D$ is an action sample for the state $s$, and $\mathbf{a}_{k}$ is the respective noisy version of $\mathbf{a}_{0}$.
Though the objectives in Eqs. (28) and (29) produce unbiased estimates of the intractable term in Eq. (11), we argue that they either lack direct interpretability—as in the case of IGL—or are overly intricate. Also, they are challenging to work with due to the exponential terms $e^{g_{\theta}}$, which can lead to training instabilities and require additional considerations during model training. For instance, CEP addressed this by self-normalizing the exponential term, while IGL mitigated it by clipping the exponential.
The proposed SWG method, conversely, is trained with the standard diffusion loss
$\mathcal{L}_{SWG}=\mathbb{E}_{k,\mathbf{z}_{0},\mathbf{z}_{k}}||{\epsilon_{0}}-{\epsilon}_{{\theta}}(\mathbf{z}_{k},k)||_{2}^{2},$ (30)
which is equivalent to maximizing a variational lower bound on the likelihood [19]. That is, a simple MSE over predicted noise terms free from exponentials prone to instabilities. Another relevant feature of SWG is that it inherently defines an in-sample loss, thus addressing the out-of-sample problem with the CEP objective.
Also, since SWG produces action and weights samples $\mathbf{z}_{0}=[\mathbf{a}_{0},w_{0}]$ simultaneously, it is possible to define a resampling mechanism as described in Sec 2.2 with the aim of further improving performance. However, our experimental results are obtained without utilizing resampling at all.
C.2 Comparison with conditional guidance
Notice that conducting conditional sampling with SWG is straightforward. Because for a label $c$ we can define the weight $w(\mathbf{a},c)\propto q(c|\mathbf{a})$, and recover the conditional sampling problem as $q(\mathbf{a}|c)\propto q(\mathbf{a})q(c|\mathbf{a})$. We claim that SWG and other exact weighted guidance methods [31, 34] generalize conditional sampling, as defining weights from label classes is straightforward—e.g., by training a classifier to estimate $q(c|\mathbf{a})$—whereas the reverse, deriving labels from weights, is not always feasible.
Appendix D Experimental details
D.1 2-D Toy examples
We followed the toy examples proposed in CEP [31], predefining a weight function $w(\mathbf{a})$ and a dataset distribution $q(\mathbf{a})$. We built the dataset $\mathcal{D}$ with 1M samples from $q(\mathbf{a})$ and extended it to $\mathcal{D}_{Z}=\{\mathbf{z}=(\mathbf{a},w(\mathbf{a}))\}_{i=1}^{N}$ using the predefined weight function.
We then trained a diffusion model $\epsilon_{\theta}$ on the extended dataset $\mathcal{D}_{Z}$. For all toy examples, the diffusion model was trained for $K=100$ diffusion steps, using a Cosine noise schedule [36]. The architecture consisted of a 6-layer Multi-Layer-Perceptron (MLP) with hidden dimension 512 and GeLU activations. We trained over $200k$ gradient steps and batch size 1024 using the Adam optimizer [20] with learning rate $41e-$ 4.
To avoid training a separate diffusion model for each $\beta$, we conditioned the diffusion model on $\beta$, $\epsilon_{\theta}(\mathbf{z}_{k},k,\beta)$, allowing us to train a single model per distribution $q(\mathbf{a})$. We trained over $\beta\in[0,20]$.
For the sampling process, we set the guidance scale $\rho$ to 1, as $\rho=1$ recovers the unbiased estimate of the target distribution score.
D.2 D4RL experiments
For the D4RL Locomotion tasks, we averaged mean returns over 5 random seeds and 20 independent evaluation trajectories per seed. For Ant Maze tasks, we averaged over 50 evaluation trajectories and 5 random seeds. The results found in Table 2 correspond to the normalized score in each domain, following the suggestion of the dataset authors [9]. Our method results correspond to the evaluation at the last training step. The results for other methods are taken from their respective original articles, except for Franka Kitchen, QGPO and IDQL results were obtained from [28] and [37], respectively.
Before training the models, we followed [9] and standardized locomotion task rewards in the dataset by dividing by the difference of returns of the best and worst trajectories in each dataset. For the Ant Maze dataset, we subtracted 1 from rewards.
To perform the offline RL experiments, it was first necessary to estimate the weight function. We used the expectile weights presented in IDQL [12] for the main D4RL results, as we found that they produce the best performance compared to the tested weights.
For the expectile weights, we needed to estimate the value functions $Q(s,\mathbf{a})$ and $V(s)$. We followed the same setup indicated in IQL [24] and IDQL [12], which consists of estimating Q-values $Q_{\psi}$ and the value function $V_{\phi}$, both with a 2-layer MLP with ReLU activations and 256 hidden units. We trained the networks with Adam optimizer [20] for 1M gradient steps, using learning rate $43e-$ 4 and batch size 256. We used double Q-learning [11] and soft updates [30] to stabilize the training. For the soft update, we used $0.005$ in all experiments to update the Q-network. We also set $\gamma=0.99$ for all tasks, and $\tau=0.7$ for the expectile loss, except for Ant Maze where we use $\tau=0.9$. The training procedure is described in Algorithm 3.
Algorithm 3 Expectile Loss critic learning
Input: Dataset $\mathcal{D}=\{(s_{i},\mathbf{a}_{i},r_{i},s^{\prime}_{i})\}_{i=1}^{N}$, discount factor $\gamma$, expectile $\tau$, batch size $B$, learning rate $\alpha$, target update parameter $\lambda$.
Initialize: Q-network $Q_{\psi}$, target Q-network $Q_{\hat{\psi}}$, value network $V_{\phi}$.
for each gradient step do
     Get $B$ samples from $\mathcal{D}$
     Compute expectile loss $L_{V}(\phi)=\mathbb{E}_{(s,\mathbf{a})\sim B}\left[L_{\tau}^{2}\left(Q_{\hat{\psi}}(s,\mathbf{a})-V_{\phi}(s)\right)\right]$
     Update value network $\phi\leftarrow\phi-\alpha\nabla_{\phi}L_{V}(\phi)$
     Compute Q-value loss $L_{Q}(\psi)=\mathbb{E}_{(r,s,\mathbf{a},s^{\prime})\sim B}\left[\left(r+\gamma V_{\phi}(s^{\prime})-Q_{\psi}(s,\mathbf{a})\right)^{2}\right]$
     Update Q-network $\psi\leftarrow\psi-\alpha\nabla_{\psi}L_{Q}(\psi)$
     Update target Q-network $\hat{\psi}\leftarrow(1-\lambda)\hat{\psi}+\lambda\psi$
end for
Return: Optimized Q-network $Q_{\psi}$ and value network $V_{\phi}$
Once the value functions are learned, we were able to calculate the weights as detailed in Sec. 4. Then, we extended our dataset to $\mathcal{D}_{Z}$ following the methodology described in Sec. 3. We trained a diffusion model $\epsilon_{\theta}$ on this extended dataset for $K=15$ diffusion steps in all experiments, using Adam optimizer [20] with learning rate $43e-$ 4 for all tasks, except for Adroit Pen where we use $53e-$ 5. We used a batch size of 1024, a variance-preserving noise schedule [50], and trained for 1M gradient steps in all tasks, except for Ant Maze, where we used 3M gradient steps.[17, 12]. The diffusion model consisted of a ResNet of 3 Residual blocks, with hidden dim 256 and Layer Normalization [2]. We used a dropout [51] of $0.1$ in all tasks. For the weight component output, we used the features from the middle layer of the ResNet and passed them through a simple MLP with a hidden dimension of 256 to obtain a more expressive representation of the weights. This architecture closely follows the one used in IDQL [12], with the exception of the additional MLP dedicated to the weight component. However, this MLP did not introduce a significant number of additional parameters. The full ResNet architecture contained approximately 1.8M parameters, ensuring a fair comparison in model capacity, as other guidance methods [31, 34] use U-Net architectures [45] with a comparable number of total parameters.
During inference, we tuned the guidance scale for each task. We swept over $\rho\in\{1,5,10,15,20,25,30\}$, such relatively large guidance scales were considered due to the expectile weights used, which were bounded in range. The guidance scale used for each experiment can be found in Tables 4, 5, 6 and 7. The results found in Table 2 correspond to the evaluation of the model at the last training step.
Table 4: Utilized guidance scales for locomotion results in Table 2
Half Cheetah Walker2d Hopper
Medium-Expert $15$ $5$ $5$
Medium $20$ $5$ $20$
Medium-Replay $30$ $10$ $10$
Table 5: Utilized guidance scales for Ant Maze results in Table 2
U-Maze Medium Large
Play/Default $5$ $15$ $10$
Diverse $1$ $15$ $10$
Table 6: Utilized guidance scales for Franka Kitchen results in Table 2
Partial Mixed Complete
$10$ $15$ $10$
Table 7: Utilized guidance scales for Adroit Pen results in Table 2
Human Cloned
$1$ $25$
D.3 Ablation over guidance scale sensitivity
We provide an ablation study on using a fixed guidance scale across all tasks, without tuning $\rho$ for each task individually, to analyze how SWG and the expectile weights were affected by the guidance scale $\rho$. The choice of the guidance scale $\rho$ controls the strength of the guidance score, which mathematically translates into a sharper target distribution concentrated around high-value weights. This effect can be seen in the modified noise prediction $\hat{\epsilon}:={\epsilon}_{\theta}(\mathbf{z}_{k},k)-\rho\sigma_{k}\nabla_{\mathbf{z}_{k}}\log{(\phi_{w}(\frac{\mathbf{z}_{k}-\sigma_{k}{\epsilon}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}}))}={\epsilon}_{\theta}(\mathbf{z}_{k},k)-\sigma_{k}\nabla_{\mathbf{z}_{k}}\log{(\phi_{w}((\frac{\mathbf{z}_{k}-\sigma_{k}{\epsilon}_{\theta}(\mathbf{z}_{k},k)}{\alpha_{k}})^{\rho}))}$.
Overall, from Tables 8 and 9, we observe that the best performance was generally achieved with $\rho=10$ or $\rho=15$. As expected, using low guidance scales ($\rho<5$) resulted in suboptimal performance, as the guidance signal was not sufficiently strong. On the other hand, employing large guidance scales ($\rho>20$) led to a drop in performance, since an overly strong guidance signal tends to produce out-of-distribution samples.
Table 8: Guidance scale $\rho$ sensitivity in Locomotion tasks.
$\rho$ 1 5 10 15 20 25 30
Average 63.8 72.0 72.9 73.1 71.2 69.2 68.8
Table 9: Guidance scale $\rho$ sensitivity in Ant Maze tasks.
$\rho$ 1 5 10 15 20 25 30
Average 24.6 37.9 69.6 68.0 67.3 61.8 57.8
D.4 Ablation over the weights experiment
To ensure a fair comparison across weights, we used the same architecture with the same training hyperparameters. To find the best performance, we also tuned over the guidance scale.
Expectile weight: The expectile weight function was learnt with the same setup adopted in the main results, described in Sec. 4.
Exponential expectile weight: For the exponential expectile weight, we used the same setup as for IDQL expectile weights. And considered a temperature $\beta=3$ to exponentiate the advantage function $e^{\beta{A(s,\mathbf{a})}}$. To ensure stability during training, we clipped the exponential weights to $(0,80]$.
Linex weight: We followed IDQL [12], and utilized $\alpha=1$ to conduct the experiments.
D.5 Scaling of inference time in SWG
Each data point in Figures 3 and 3 represents the average inference time with our method for $K=15$ diffusion steps, obtained over 10,000 runs of the algorithm, along with the corresponding error region. For completeness, we report the total number of parameters in each network used in our experiments. The table below shows the parameter counts corresponding to different width and depth configurations:
Table 10: Number of parameters for each network configuration used in our experiments. Left: scaling with width (hidden size). Right: scaling with depth (number of layers).
Width (hidden size) Parameters
128 180K
256 370K
512 1M
1024 3.5M
2056 13.3M
4096 51.5M
8192 203.7M
Depth (layers) Parameters
1 660K
2 4.5M
3 9.1M
4 13.3M
5 17.5M
6 21.8M
7 26M
D.6 Computational resources
Our implementation is based on the jaxrl repository [23] using the JAX [3] and Flax [13] libraries. We also provide an implementation of our method in PyTorch [38]. All experiments were conducted on a 12GB NVIDIA RTX 3080 Ti GPU. Training takes approximately 0.5 hours for value function learning (i.e., weight training, 1M steps), and an additional 0.5 hours for 1M steps of the diffusion model. However, since experiments can be run in parallel, the overall training time can be significantly reduced. The code used for training and experimentation can be found in SWG repository
Appendix E More toy examples
Figure 4: Additional toy examples with our proposed method – see Sec. 5.1 for description.